3 + sqrt 3x - 2 = -5
Answer by jerrybassmasta(14)

3 + sqrt 3x - 2 = -5
To solve this is pretty simple if you get the square root.
First ask yourself, what is the opposite of a square root?
If you know that you can solve the problem easily.
You have to square it.
3 + sqrt 3x - 2 = -5
becomes
3+3x-2=-5^2
Following PEMDAS, you have to solve the exponent first.
3+3x-2=25
Add 2 on the other side and then subtract 3.
3x=24
Now divide.
x=8